OpenStudy (anonymous):

A worker in an assembly line takes 2 hours to produce 120 parts. At that rate how many parts can she produce in 9 hours?

jimthompson5910 (jim_thompson5910):

2/120 = 9/x solve for x

OpenStudy (anonymous):

i dont get it

jimthompson5910 (jim_thompson5910):

cross multiply to get 2/120 = 9/x 2*x = 120*9 2x = 1080 x = ???

OpenStudy (anonymous):

540

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (whpalmer4):

In general, find the unit rate: amount done in 1 unit of time. Then multiply by the amount of time. Here we have 120 parts in 2 hours, so the unit rate is 120 parts / 2 hours = 60 parts/hour. Now multiply that by 9 hours and you get 60 parts / hour * 9 hours = 540 parts.